OpenStudy (anonymous):

Limits lim_{x rightarrow infty } sqrt{x ^{2} + x} - sqrt{x ^{2} + 1}

OpenStudy (bahrom7893):

http://www.wolframalpha.com/input/?i=lim+%28sqrt%28x%5E2%2Bx%29-sqrt%28x%5E2%2B1%29%2Cx%2Cinfinity%29 Click on show steps to see the steps.. Wolfram is your best friend (for the most problems)...

OpenStudy (anonymous):

I do not understand with this form

OpenStudy (bahrom7893):

what form? indeterminate form?

hartnn (hartnn):

can u use L'Hopitals ?

OpenStudy (anonymous):

no @hartnn

OpenStudy (anonymous):

yes @bahrom7893

OpenStudy (bahrom7893):

Ohhhhhh i see

hartnn (hartnn):

try this, put x=1/y as x->infinity, y->0

hartnn (hartnn):

u will get : \(\large \frac{\sqrt{1+y}-\sqrt{1+y^2}}{y}\) now multiply by conjugate.

hartnn (hartnn):

are u trying ? @Muskan or should i give further steps ?

OpenStudy (anonymous):

give me setps

hartnn (hartnn):

first, did u get my substitution ? and how i got that function in y ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

\(\large \frac{\sqrt{1+y}-\sqrt{1+y^2}}{y}\quad \large \frac{\sqrt{1+y} +\sqrt{1+y^2}}{\sqrt{1+y} +\sqrt{1+y^2}}=\frac{1+y-1+y^2}{y*(\sqrt{1+y} +\sqrt{1+y^2})}\) can u solve further ?

OpenStudy (anonymous):